What is the value of x in the equation
(2x+2)+(3x-12)=180 ?

Answer by Cromlix(4381)

Hi there,
(2x+2)+(3x-12)=180
Remove brackets.
2x + 2 + 3x - 12 = 180
Collect like terms:
2x + 3x = 180 - 2 + 12
5x = 190
x = 38
Hope this helps :-)